OpenStudy (anonymous):

The sum of Pete's and Sam's ages is 30. Five years ago, Pete was 3 times as old as Sam. How old is Sam? Let P = Pete's age, S = Sam's age, and P + S = 30. Which of the following equations would complete the system? P - 5 = 3S P - 5 = 3S - 5 P - 5 = 3S - 15

OpenStudy (anonymous):

A=B=C=D=E THEN FMUST = 1

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

wat

OpenStudy (wolfe8):

So First statement: The sum of Pete's and Sam's ages is 30 P+S=30 Second statement: Five years ago, Pete was 3 times as old as Sam P-5+S-5=P+3S

OpenStudy (anonymous):

go on

OpenStudy (wolfe8):

That's it. You are asked what statement would complete the system, which is the second statement.

OpenStudy (anonymous):

oh lol thanks dood/chick